# Inference for Deep Neural Network Estimators in Generalized Nonparametric Models

### v3

@@ 3 Theoretical Results
Suppose that (2.3) holds and the true function $f_{0}\in\mathcal{G}(q,\mathbf{d},\mathbf{t},{\bm{\gamma}},K)$. The support of $y$, denoted by $\mathrm{supp}(h(y))$, is fixed and does not depend on $\mathbf{x}$. Moreover, the sufficient statistic $T(y)=y$ is not degenerate and can take at least two distinct values in $\mathrm{supp}(h)$. For any $\mathbf{x}$, the conditional distribution $p(y|\mathbf{x})$ is sub-exponential. Specifically, with bounded $f_{0}$ there exists a universal constant $\kappa>0$ such that for all $t>0$,
@@
The network class $\mathcal{F}(L,\mathbf{p},s,F)$ with $L$, $\mathbf{p}$ and $s$ satisfies i) $F\geq\max(K,1)$, ii) $\sum_{i=0}^{q}\log_{2}(4t_{i}\vee 4\gamma_{i})\log_{2}n\leq L\precsim\log^{\alpha}n$ for some $\alpha>1$, iii) $n\phi_{n}\precsim\min_{i=1,...,L}p_{i}\leq\max_{i=1,...,L}p_{i}\precsim n$, iv) $s\precsim n\phi_{n}\log n$. Here, $\phi_{n}=\max_{j=0,...,q}n^{-2\gamma_{j}^{*}/(2\gamma_{j}^{*}+t_{j})}$, where $\gamma_{j}^{*}=\gamma_{j}\Pi_{l=i+1}^{q}(\gamma_{l}\wedge 1)$. The values $t_{i},\gamma_{i}$ and $p_{i}$ are the $(i+1)$-th element in the vectors $\mathbf{t}$, $\bm{\gamma}$} and $\mathbf{p}$, respectively.
@@
The covariance $\xi_{1,r}(\mathbf{x})$ defined in (3.1) satisfies $\inf_{\mathbf{x}\in\mathcal{X}}\xi_{1,r}(\mathbf{x})\geq Cr^{-(1+\varepsilon)}$ for some small $\varepsilon>0$. The subsample size $r=n^{\gamma}$ satisfies $1/(1+\min_{j=0,...,q}(2\gamma_{j}^{*})/(2\gamma_{j}^{*}+t_{j}))<\gamma<1$, and the number of subsampling times $B$ is large such that $B\succsim n$.
Assumption 3.1 ensures that $p(y|\mathbf{x})$ not degenerate, while the concentration of $y$ does not deviate too widely. Additionally, that $T(y)=y$ takes at least two distinct values in $\mathrm{supp}(h)$ guarantees a non-zero variance of $y$, ensuring the log-partition function $\psi(f_{0}(\mathbf{x}))$ exhibits local convexity. The boundedness of $f_{0}$ guarantees the existence of $\kappa$. This condition, weaker than the sub-Gaussian tail bound assumed in Fan et al. (2024), is satisfied by most common exponential family distributions, including Beta, Bernoulli, Poisson, and Gaussian. Assumption 3.2 is a technical assumption similar to Schmidt-Hieber (2020), which specifies the requirements on the network parameters relative to the sample size $n$ and smoothness parameters $\phi_{n}$, $t_{i}$ and $\gamma_{i}$, including the number of layers $L$ and the sparsity $s$. Assumption 3.3, related to the training dynamics of DNNs, stipulates that DNNs return an estimator sufficiently close to the global minimum. We note that the rate in Assumption 3.3 pertains to the prediction error in Theorem 3.5. Our intent is not to assert an explicit convergence rate for SGD in practice, but to impose a sufficient condition that separates optimization error from the inferential analysis. This perspective follows standard practice in the statistical literature (Chernozhukov et al., 2018; Schmidt-Hieber, 2020; Fan & Gu, 2024), where such assumptions serve as technical devices to ensure valid inference rather than as primary objects of study. Assumption 3.4 on $\xi_{1,r}(\mathbf{x})$ is not meant to serve as a precise characterization of neural networks. Rather, it ensures that the first-order Hoeffding projection remains non-degenerate and dominates the higher-order components, as in the random forest setting (Wager & Athey, 2018). We choose the subsample size $r$ much smaller than $n$ to ensure first-order dominance but large enough to limit bias, and let the number of subsamples $B$ scale with $n$ to control computation.
@@
Suppose that $f_{0}\in\mathcal{G}(q,\mathbf{d},\mathbf{t},{\bm{\gamma}},K)$ and $\widehat{f}_{n}\in\mathcal{F}(L,\mathbf{p},s,F)$. If Assumptions 3.1-3.2 hold, then it holds that
@@
Theorem 3.5 establishes a bound on the out-of-sample estimation error. Unlike previous studies (Schmidt-Hieber, 2020; Fan & Gu, 2024), which primarily focus on bounding the in-sample mean squared error, i.e., $\frac{1}{n}\sum_{i=1}^{n}(y_{i}-f(\mathbf{x}_{i}))^{2}$, to derive bounds for the estimation error, our approach applies to a wide class of training losses (2.6) under GNRMs. The heteroskedasticity inherent in GNRMs necessitates a more refined analysis, and a truncation technique is introduced to address this challenge by separating the noise into different regimes. Specifically, this result enables the analysis of non-Gaussian outcomes with heteroskedasticity depending on covariates, overcoming key limitations of prior methods (Schmidt-Hieber, 2020; Fan & Gu, 2024) that rely on the independence of $y_{i}-\mathbb{E}(y_{i}|\mathbf{x}_{i})$ and $f_{0}(\mathbf{x}_{i})$ to derive estimation error bounds. Theorem 3.5 also implies that the aggregated estimator $\widehat{f}^{B}$ has a vanishing bias, as shown in the following corollary.
Corollary 3.6.
Assume that the conditions of Theorem 3.5 hold, the neural network fitted on subsamples belongs to $\mathcal{F}(L,\mathbf{p},s,F)$, Assumptions 3.2 and 3.3 hold for each subsample with size $r$, and Assumption 3.4 holds. Then for a new independent and identical random variable $\bm{X}\sim\mathbb{P}_{\mathbf{x}}$, it holds that
$\frac{n}{r^{2}\inf_{\mathbf{x}\in\mathcal{X}}\xi_{1,r}(\mathbf{x})}\cdot\mathbb{E}\big|\mathbb{E}(\widehat{f}^{B}(\bm{X})-f_{0}(\bm{X})|\bm{X})\big|^{2}~\to~0.$
Corollary 3.6 implies asymptotic unbiasedness, i.e., for a fixed $\mathbf{x}_{*}$, $\mathbb{E}\{\widehat{f}^{B}(\mathbf{x}_{*})\}-f_{0}(\mathbf{x}_{*})=o\Big(\sqrt{\frac{r^{2}\xi_{1,r}(\mathbf{x}_{*})}{n}}\Big)$. Thus, the bias of $\widehat{f}^{B}$ is asymptotically negligible, allowing us to focus on $\widehat{f}^{B}(\mathbf{x}_{*})-\mathbb{E}\{\widehat{f}^{B}(\mathbf{x}_{*})\}$ and establish the following results.
@@
Suppose that the data $\mathcal{D}_{n}$ are generated from Model (2.3) with an unknown $f_{0}\in\mathcal{G}(q,\mathbf{d},\mathbf{t},{\bm{\gamma}},K)$, the used neural networks belong to $\mathcal{F}(L,\mathbf{p},s,F)$, Assumptions 3.1, 3.3 and 3.4 hold, and Assumption 3.2 hold for each subsample with size $r$. Then there exists a positive sequence, $\delta_{1},\delta_{2},\ldots,$ with $\delta_{n}\to 0$ and corresponding set $\mathcal{A}_{\delta_{n}}\subseteq\mathcal{X}$ with $\mathbb{P}_{\mathbf{x}}(\mathcal{A}_{\delta_{n}})\geq 1-\delta_{n}$, such that for any fixed $\mathbf{x}_{*}\in\mathcal{A}_{\delta_{n}}$, the estimator $\widehat{f}^{B}(\mathbf{x}_{*})$ satisfies
@@
Compared to Wager & Athey (2018) and Fei & Li (2024), since we allow the number of subsamples $B$ to grow at the same rate as the sample size $n$, the Monte Carlo bias in the variance estimation becomes non-negligible. As pointed out by Wager et al. (2014), this bias is typically of order $n/B$, and does not vanish when $B\asymp n$. Different from the correction proposed in Wager et al. (2014), which assumes near asymptotic independence between the inclusion indicators $J_{bi}$ and the estimates $\widehat{f}^{b}$, we propose a bias-correction term that directly accounts for the dependency between subsample structure and estimates. Specifically, we utilize the covariance form of the variance estimator and subtract a U-statistic–based empirical correction that estimates the within-subject variability due to repeated subsampling. This yields a debiased estimator that is asymptotically unbiased even when $B\asymp n$. For the confidence interval in (2.8), Theorem 3.8 and the Slutsky theorem ensure that, as $n\to\infty$, it holds that
@@
Remark. Our inference framework relies on a well-trained network; when training is insufficient or overly regularized, inference may become unreliable. For analytical tractability, we adopt an idealized $s$-sparse, $F$-bounded network class (Schmidt-Hieber, 2020), while in practice, regularization techniques such as weight decay and dropout promote approximate sparsity, keeping the trained model close to this class. The optimization gap $\Delta_{n}(\widehat{f}_{n})$ in Theorem 3.5 quantifies the discrepancy between empirical and population risks, filters out degenerate training cases, and accounts for residual optimization and approximation errors. Our theory demonstrates that as long as $\Delta_{n}(\widehat{f}_{n})$ remains small, inference validity is preserved, consistent with Chernozhukov et al. (2018); Fan & Gu (2024).
@@ 4 Numerical Experiments
We present the comparison results in Table 2. For ESM, we select our results with $\gamma=0.9$. For HulC method proposed by Kuchibhotla et al. (2024), although the coverage probability is close to the nominal level of $95\%$, its average interval length is very wide. For Bayesian neural networks (Sun et al., 2021), we observe that the performance is highly sensitive to the tuning of hyperparameters. Due to the difficulty of selecting appropriate priors and optimization settings, the coverage probabilities and average interval lengths in our experiments are worse than those of the proposed ESM, with noticeably wider intervals and less stable coverage. For the ensemble methods (Lakshminarayanan et al., 2017) and the naive bootstrap approach, the intervals are wider than those of ESM. As the sample size $n$ increases, and the coverage probability tends to exceed the nominal $95\%$. In comparison, our ESM consistently achieves near-nominal coverage with shorter interval lengths, providing higher efficiency under accuracy guarantees.
@@
We also investigate the performance of narrower yet deeper architectures under our method. We design a deeper network with $L=5$ and $\mathbf{p}=(p,10,15,20,30,1)$. As deep networks may cause gradient explosion or vanishing during backpropagation (Glorot & Bengio, 2010), we incorporate batch normalization layers after the second and second-to-last layers, aiming to mitigate gradient instability while maintaining the network’s ability to learn complex representations. Table 3 shows that key metrics, including empSD, SE, and $\text{SE}_{c}$, remain largely consistent with those observed for shallower network architectures. Nevertheless, coverage probabilities drop to around 91%. Figure 3 presents representative results for $n=400$ and $r=n^{0.8}$. As shown in Figure 3(a), the fitted values exhibit a slight systematic deviation from the diagonal, indicating bias relative to the true function $f_{0}$. Nevertheless, Figure 3(b) confirms that the corrected variance estimates remain accurate. This bias shifts the confidence intervals, leading to reduced coverage despite precise variance estimation. Similar bias-related undercoverage has been reported for random forests in Wager & Athey (2018). In our case, greater network depth may induce complex interactions that, without sufficient data, amplify estimation bias.
@@ 6 Discussion
[…]
Our theoretical analysis focuses on ReLU activations, where monomial-based approximations are tractable. We envision that ESM can be extended to networks with smooth activation functions such as tanh or GELU, although this requires careful consideration. Unlike ReLU, smooth activations are less flexible in local adaptations, which may result in slower approximation rates and influence convergence behavior. To preserve inference validity in this setting, it may be necessary to employ alternative analytical tools or apply stronger regularization to control model complexity and prevent saturation. Moreover, our analysis assumes that DNN estimators, trained via gradient descent, effectively minimize empirical risk and generalize well, an increasingly supported view in recent theory. While we do not analyze training dynamics explicitly, our assumptions are consistent with emerging work on implicit bias and convergence in overparameterized models (Chizat & Bach, 2020; Chatterji et al., 2021; Cao et al., 2022). An important direction is to examine how implicit regularization techniques, such as early stopping, dropout, and weight decay, enhance ESM’s applicability by promoting model stability. These methods curb overfitting, reduce sensitivity to data perturbations, and help satisfy the smoothness and exchangeability conditions required for valid U-statistic–based inference.

### v4

@@ 3 Theoretical Results
Suppose that (2.3) holds and the true function $f_{0}\in\mathcal{G}(q,\mathbf{d},\mathbf{t},\mathbf{m},K)$. The support of $y$, denoted by $\mathrm{supp}(h(y))$, is fixed and does not depend on $\mathbf{x}$. Moreover, the sufficient statistic $T(y)=y$ is not degenerate and can take at least two distinct values in $\mathrm{supp}(h)$. For any $\mathbf{x}$, the conditional distribution $p(y|\mathbf{x})$ is sub-exponential. Specifically, with bounded $f_{0}$ there exists a universal constant $\kappa>0$ such that for all $t>0$,
@@
The network class $\mathcal{F}(L,\mathbf{p},s,F)$ with $L$, $\mathbf{p}$ and $s$ satisfies i) $F\geq\max(K,1)$, ii) $\sum_{i=0}^{q}\log_{2}(4t_{i}\vee 4m_{i})\log_{2}n\leq L\precsim\log^{\alpha}n$ for some $\alpha>1$, iii) $n\phi_{n}\precsim\min_{i=1,...,L}p_{i}\leq\max_{i=1,...,L}p_{i}\precsim n$, iv) $s\precsim n\phi_{n}\log n$. Here, $\phi_{n}=\max_{j=0,...,q}n^{-2m_{j}^{*}/(2m_{j}^{*}+t_{j})}$, where $m_{j}^{*}=m_{j}\Pi_{l=i+1}^{q}(m_{l}\wedge 1)$. The values $t_{i},m_{i}$ and $p_{i}$ are the $(i+1)$-th element in the vectors $\mathbf{t}$, $\mathbf{m}$ and $\mathbf{p}$, respectively.
@@
The covariance $\xi_{1,r}(\mathbf{x})$ defined in (3.1) satisfies $\inf_{\mathbf{x}\in\mathcal{X}}\xi_{1,r}(\mathbf{x})\geq Cr^{-(1+\varepsilon)}$ for some $\varepsilon>0$. The subsample size $r=n^{\gamma}$ satisfies $1/(1+\min_{j=0,...,q}(2m_{j}^{*})/(2m_{j}^{*}+t_{j}))<\gamma<1$, and the number of subsampling times $B$ is large such that $B\succsim n$.
Assumption 3.1 ensures that $p(y|\mathbf{x})$ not degenerate, while the concentration of $y$ does not deviate too widely. Additionally, that $T(y)=y$ takes at least two distinct values in $\mathrm{supp}(h)$ guarantees a non-zero variance of $y$, ensuring the log-partition function $\psi(f_{0}(\mathbf{x}))$ exhibits local convexity. The boundedness of $f_{0}$ guarantees the existence of $\kappa$. This condition, weaker than the sub-Gaussian tail bound assumed in Fan et al. (2024), is satisfied by most common exponential family distributions, including Beta, Bernoulli, Poisson, and Gaussian. Assumption 3.2 is a technical assumption similar to Schmidt-Hieber (2020), which specifies the requirements on the network parameters relative to the sample size $n$ and smoothness parameters $\phi_{n}$, $t_{i}$ and $m_{i}$, including the number of layers $L$ and the sparsity $s$. Assumption 3.3, related to the training dynamics of DNNs, stipulates that DNNs return an estimator sufficiently close to the global minimum. We note that the rate in Assumption 3.3 pertains to the prediction error in Theorem 3.5. Our intent is not to assert an explicit convergence rate for SGD in practice, but to impose a sufficient condition that separates optimization error from the inferential analysis. This perspective follows standard practice in the statistical literature (Chernozhukov et al., 2018; Schmidt-Hieber, 2020; Fan and Gu, 2024), where such assumptions serve as technical devices to ensure valid inference rather than as primary objects of study. Assumption 3.4 on $\xi_{1,r}(\mathbf{x})$ is not meant to serve as a precise characterization of neural networks. Rather, it ensures that the first-order Hoeffding projection remains non-degenerate and dominates the higher-order components, as in the random forest setting (Wager and Athey, 2018). We choose the subsample size $r$ much smaller than $n$ to ensure first-order dominance but large enough to limit bias, and let the number of subsamples $B$ scale with $n$ to control computation.
@@
Suppose that $f_{0}\in\mathcal{G}(q,\mathbf{d},\mathbf{t},\mathbf{m},K)$ and $\widehat{f}_{n}\in\mathcal{F}(L,\mathbf{p},s,F)$. If Assumptions 3.1-3.2 hold, then it holds that
@@
Theorem 3.5 establishes a bound on the out-of-sample estimation error. Unlike previous studies (Schmidt-Hieber, 2020; Fan and Gu, 2024), which primarily focus on bounding the in-sample mean squared error, i.e., $\frac{1}{n}\sum_{i=1}^{n}(y_{i}-f(\mathbf{x}_{i}))^{2}$, to derive bounds for the estimation error, our approach applies to a wide class of training losses (2.6) under GNRMs. The heteroskedasticity inherent in GNRMs necessitates a more refined analysis, and a truncation technique is introduced to address this challenge by separating the noise into different regimes. Specifically, this result enables the analysis of non-Gaussian outcomes with heteroskedasticity depending on covariates, overcoming key limitations of prior methods (Schmidt-Hieber, 2020; Fan and Gu, 2024) that rely on the independence of $y_{i}-\mathbb{E}(y_{i}|\mathbf{x}_{i})$ and $f_{0}(\mathbf{x}_{i})$ to get estimation error bounds.
Proposition 3.6.
We suppose that the conditions of Theorem 3.5 hold, the neural network fitted on subsamples belongs to $\mathcal{F}(L,\mathbf{p},s,F)$, Assumptions 3.2-3.4 hold for each subsample with size $r$, and sparsity $s\asymp r^{1+\delta}\phi_{r}\log r$ for some constant $\delta>0$. Then for a fixed $\mathbf{x}_{*}$, if the pointwise bias is dominated by the best approximation bias, i.e., there exists constant $C(\mathbf{x}_{*})$ only depends on $\mathbf{x}_{*}$ such that $\big|\mathbb{E}\widehat{f}^{b}(\mathbf{x}_{*})-f_{0}(\mathbf{x}_{*})\big|\precsim C(\mathbf{x}_{*})\inf_{f\in\mathcal{F}}\bigl|f(\mathbf{x}_{*})-f_{0}(\mathbf{x}_{*})\bigr|$, it holds that $\sqrt{\frac{n}{r^{2}\inf_{\mathbf{x}\in\mathcal{X}}\xi_{1,r}(\mathbf{x})}}\cdot\big|\mathbb{E}\widehat{f}^{B}(\mathbf{x}_{*})-f_{0}(\mathbf{x}_{*})\big|~\to~0$.
Proposition 3.6 implies asymptotic unbiasedness, i.e., $\mathbb{E}\{\widehat{f}^{B}(\mathbf{x}_{*})\}-f_{0}(\mathbf{x}_{*})=o\Big(\sqrt{\frac{r^{2}\xi_{1,r}(\mathbf{x}_{*})}{n}}\Big)$ for a fixed $\mathbf{x}_{*}$. Thus, the bias of $\widehat{f}^{B}$ is asymptotically negligible, and enables us to focus on $\widehat{f}^{B}(\mathbf{x}_{*})-\mathbb{E}\{\widehat{f}^{B}(\mathbf{x}_{*})\}$ and establish the following results.
@@
Suppose the data $\mathcal{D}_{n}$ are generated from Model (2.3) with an unknown $f_{0}\in\mathcal{G}(q,\mathbf{d},\mathbf{t},\mathbf{m},K)$, the used neural networks belong to $\mathcal{F}(L,\mathbf{p},s,F)$, and Assumptions 3.13.4 hold (Assumption 3.2 holds for each subsample with size $r$). Then with the results of Proposition 3.6, there exists a positive sequence, $\delta_{1},\delta_{2},\ldots,$ with $\delta_{n}\to 0$ and a corresponding set $\mathcal{A}_{\delta_{n}}\subseteq\mathcal{X}$ with $\mathbb{P}_{\mathbf{x}}(\mathcal{A}_{\delta_{n}})\geq 1-\delta_{n}$, such that for any fixed $\mathbf{x}_{*}\in\mathcal{A}_{\delta_{n}}$, it holds that
@@
Compared to Wager and Athey (2018) and Fei and Li (2024), since we allow the number of subsamples $B$ to grow at the same rate as the sample size $n$, the Monte Carlo bias in the variance estimation becomes non-negligible. As pointed out by Wager et al. (2014), this bias is typically of order $n/B$, and does not vanish when $B\asymp n$. Different from the correction proposed in Wager et al. (2014), which assumes near asymptotic independence between the inclusion indicators $J_{bi}$ and the estimates $\widehat{f}^{b}$, we propose a bias-correction term that directly accounts for the dependency between subsample structure and estimates. Specifically, we utilize the covariance form of the variance estimator and subtract a U-statistic–based empirical correction that estimates the within-subject variability due to repeated subsampling. This yields a debiased estimator that is asymptotically unbiased even when $B\asymp n$. For the confidence interval in (2.8), Theorem 3.8 and the Slutsky theorem ensure that, as $n\to\infty$, it holds that
@@
Remark. Our inference framework relies on a well-trained network; when training is insufficient or overly regularized, inference may become unreliable. For analytical tractability, we adopt an idealized $s$-sparse, $F$-bounded network class (Schmidt-Hieber, 2020), while in practice, regularization techniques such as weight decay and dropout promote approximate sparsity, keeping the trained model close to this class. The optimization gap $\Delta_{n}(\widehat{f}_{n})$ in Theorem 3.5 quantifies the discrepancy between empirical and population risks, filters out degenerate training cases, and accounts for residual optimization and approximation errors. Our theory demonstrates that as long as $\Delta_{n}(\widehat{f}_{n})$ remains small, inference validity is preserved, consistent with Chernozhukov et al. (2018); Fan and Gu (2024).
@@ 4 Numerical Experiments
We present the comparison results in Table 2. For ESM, we select our results with $\gamma=0.9$. For HulC method proposed by Kuchibhotla et al. (2024), although the coverage probability is close to the nominal level of $95\%$, its average interval length is very wide. For Bayesian neural networks (Sun et al., 2021), we observe that the performance is highly sensitive to the tuning of hyperparameters. Due to the difficulty of selecting appropriate priors and optimization settings, the coverage probabilities and average interval lengths in our experiments are worse than those of the proposed ESM, with noticeably wider intervals and less stable coverage. For the ensemble methods (Lakshminarayanan et al., 2017) and the naive bootstrap approach, the intervals are wider than those of ESM. As the sample size $n$ increases, and the coverage probability tends to exceed the nominal $95\%$. Overall, ESM achieves competitive coverage performance with narrowest intervals, although it may not always provide coverage closest to the nominal level among the compared methods. So, in some cases, using estimators without bias correction or using bootstrap procedures may also have merit, as the presence of bias-correction terms in finite samples can lead to slight undercoverage.
@@
We also investigate the performance of narrower yet deeper architectures under our method. We design a deeper network with $L=5$ and $\mathbf{p}=(p,10,15,20,30,1)$. As deep networks may cause gradient explosion or vanishing during backpropagation (Glorot and Bengio, 2010), we incorporate batch normalization layers after the second and second-to-last layers, aiming to mitigate gradient instability while maintaining the network’s ability to learn complex representations. Table 3 shows that key metrics, including empSD, SE, and $\text{SE}_{c}$, remain largely consistent with those observed for shallower network architectures. Nevertheless, coverage probabilities drop to around 91%. Figure 3 presents representative results for $n=400$ and $r=n^{0.8}$. As shown in Figure 3(a), the fitted values exhibit a slight systematic deviation from the diagonal, indicating bias relative to the true function $f_{0}$. Nevertheless, Figure 3(b) confirms that the corrected variance estimates remain accurate. This bias shifts the confidence intervals, leading to reduced coverage despite precise variance estimation. Similar bias-related undercoverage has been reported for random forests in Wager and Athey (2018). In our case, greater network depth may induce complex interactions that, without sufficient data, amplify estimation bias.
@@ 6 Discussion
Our methodology may have broad applicability as it enables valid inference for any continuously-differentiable functional of $f_{0}(x)$, such as $\operatorname{{\rm Var}}(y|\mathbf{x})=\psi^{\prime\prime}(f_{0}(x))=:g(f_{0}(x))$, where $g(u)=\psi^{\prime\prime}(u)$, the second derivative of $\psi(u)$. We can estimate $\operatorname{{\rm Var}}(y|\mathbf{x})$ via the plug-in estimator $g(\widehat{f}^{B}(\mathbf{x}))$ and apply the delta method to give $\sqrt{\frac{n}{r^{2}\xi_{1,r}(\mathbf{x})}}\cdot(g(\widehat{f}^{B}(\mathbf{x}))-g(f_{0}(\mathbf{x})))=\sqrt{\frac{n}{r^{2}\xi_{1,r}(\mathbf{x})}}\cdot g^{\prime}(f_{0}(\mathbf{x}))(\widehat{f}^{B}(\mathbf{x})-f_{0}(\mathbf{x}))+o_{p}(1)$, which will lead to estimated confidence intervals. Our method may also naturally extend to semiparametric causal inference settings where the target parameter can be expressed as a functional of the conditional means. Specifically, to estimate the Conditional Average Treatment Effect (CATE), $\tau(\mathbf{x})=\mathbb{E}[y(1)|\mathbf{x}]-\mathbb{E}[y(0)|\mathbf{x}]$, where $y(1)$ and $y(0)$ denote the potential outcomes under treatment and control, we apply ESM separately to treated and control groups under standard unconfoundedness. Multiple subsamples are drawn for each group, predictive models are trained to estimate $\mathbb{E}[y(g)|\mathbf{x}]$, $g=0,1$, and ensemble means $\widehat{\mu}_{1}(\mathbf{x})$ and $\widehat{\mu}_{0}(\mathbf{x})$ are obtained. The CATE estimator is $\widehat{\tau}(\mathbf{x})=\widehat{\mu}_{1}(\mathbf{x})-\widehat{\mu}_{0}(\mathbf{x}).$ ESM’s variance framework provides inference by obtaining $\widehat{\sigma}_{g}^{2}(\mathbf{x})=\widehat{\operatorname{{\rm Var}}}\{\widehat{\mu}_{g}(\mathbf{x})\}$ for $g=0,1$, yielding pointwise confidence intervals $\widehat{\tau}(\mathbf{x})\pm z_{1-\alpha/2}\sqrt{\widehat{\sigma}_{1}^{2}(\mathbf{x})+\widehat{\sigma}_{0}^{2}(\mathbf{x})}.$ Detailed justifications warrant investigation.
Our theoretical analysis focuses on ReLU activations, where monomial-based approximations are tractable. We envision that ESM can be extended to networks with smooth activation functions such as tanh or GELU, although this requires careful consideration. Unlike ReLU, smooth activations are less flexible in local adaptations, which may result in slower approximation rates and influence convergence behavior. To preserve inference validity in this setting, it may be necessary to employ alternative analytical tools or apply stronger regularization to control model complexity and prevent saturation. Moreover, Our theoretical analysis relies on several assumptions that are broadly consistent with recent work on implicit bias and convergence in overparameterized models (Chizat and Bach, 2020; Chatterji et al., 2021; Cao et al., 2022). We acknowledge, however, that some of these assumptions may appear strong. In particular, the conditions imposed on $\xi_{1,r}(\mathbf{x})$, which ensure the validity of the first-order Hoeffding projection, can be further improved. An important direction for future work is to relax these conditions by linking them more directly to of the neural network class, such as its sparsity or other structural characteristics.
@@
Acknowledgements. We are deeply grateful to the Editor, Associate Editor, two referees and one reproducibility reviewer for insightful reviews that have greatly improved the manuscript. This work was supported by NIH grants 2R01CA249096-5 and R01CA269398.